Sales are planned for a department at $2,345,698 for the Fall season. The seasonal turnover objective is 1.6.  Calculate the BOM stock for the month of September if sales are planned at $324,643 and the basic stock method is used.

$875,727

$1,274,495

$987,333

$2,345,698

None of the above
Answer #1

We would use the following formulas:

Basic Stock= Average Stock for the Season - Average Monthly Sales for the Season

where

  Average Monthly Sales for the Season = Total Planned Sales for the Season / Number of months in the Season

Average Stock for the Season = Total Planned Sales for the Season / Estimated Inventory Turnover Rate for the Season

Beginning of the Month (BOM) stock = Planned Monthly Sales + Basic Stock

Fall Season is generally for 1 months:

  thus Average Monthly Sales for the Season = 2345698/1= 2345698   
Average Stock for the Season = 2345698/1.6 = 1466061.25  

thus Basic Stock = 1466061.25 - 2345698 =-879636.75

Thus September BOM Stock : Planned Monthly Sales (324,643) + Basic Stock (-879636.75 rounding off)          
: 324,643+(- 879636.75)= -554993.75

ANS: None of the above